OpenStudy (anonymous):

How do you find all values of s that satisfy the Mean Value Theorem for f(s)=(4s+16)^(2/3) on the interval [-7,-4] without a calculator?

OpenStudy (anonymous):

Mean Value Theorem guarantees a value of c on [-7,-4] such that f'(c) = [f(-4)-f(-7)]/(-4 - (-7)] So you need the derivative of the function.

OpenStudy (anonymous):

I did that, but its determining the correct way to solve it that I have trouble with. I found the derivative as \[\frac{ 8 }{ 3\sqrt[3]{4s+16} }\] and set it equal to the average slope, \[\frac{ -12^{2/3} }{ -3 }\], but I don't know what do after that.

myininaya (myininaya):

Hey. I don't like my s being on bottom. Maybe you might want to multiply both sides by that cube root part or even multiply both sides by 3 time that cube root part.

myininaya (myininaya):

Or if the thing that makes you mad is the cube root, then cube both sides to get rid of the cube root.

myininaya (myininaya):

I would end up doing the combination of the two things. :)